# Trees of Graphs as Boundaries of Hyperbolic Groups

### v1

@@ 13.1. Description of $\Theta^{V}[{\mathcal{R}}_{G}]$ the tree system of V-trees associated to $G$
      moreover, if $A$ is ##\##-non-orientable and $\varepsilon$ is orientable then, after fixing an auxiliary orientation of $A$ and some orientation for $\varepsilon$, the union of all those E-peripherals $P$ of $A$ for which the adjacent to $P$ peripheral $P^{\prime}$ is contained in an arm or wedge arm $A^{\prime}$ of U-type $\varepsilon$, and for which the orientation induced from the fixed orientation of $\varepsilon$ (via the induced orientations of $A^{\prime}$ and $P^{\prime}$) is compatible (respectively, incompatible) with the fixed orientation of $A$, is dense in $A$.
We finish this subsection with another two observations concerning the just introduced concepts, which describe some features of the C-peripherals of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$. These observations follow rather easily from Facts 10.2 and 10.3, respectively, and we omit their proofs.
@@
    For all $s\in N_{u}^{T_{*}}\setminus\{t\}$ for which the groups $G(b(s))$ are orientable (as flexible factors), the arms $K^{*}_{s}$ are ##\##-orientable, and the ##\##-orientations of these arms are compatible, i.e. they all induce (via the maps $(\varphi_{[s,u]}^{*})^{-1}$) the same order on the doubleton $\Sigma^{*}_{u}$.
  4. (3)
    For any $s$ and $\varepsilon$ as in (1), if the group $G(b(s))$ is non-orientable (as a flexible factor), and if the edge group $G([u_{\xi},b(s)])$ is orientable, then the cardinality of the set of arms $K^{*}_{s^{\prime}}$ with $s^{\prime}\in N_{u}^{T_{*}}\setminus\{t\}$ and with U-type $\varepsilon$ is even, all these arms are ##\##-orientable, and the ##\##-orientations of half of these arms are compatible with one of the orders of the doubleton $\Sigma^{*}_{u}$, while the ##\##-orientations of the other half of these arms are compatible with the opposite order of $\Sigma^{*}_{u}$.
@@
    For all $s\in N_{u}^{T_{*}}$ for which the groups $G(b(s))$ are orientable (as flexible factors), the wedge arms $K^{*}_{s}$ are ##\##-orientable, and the ##\##-orientations of these arms are compatible, i.e. they all induce (via the maps $(\varphi_{[s,u]}^{*})^{-1}$) the same order on the doubleton $\Sigma^{*}_{u}$.
  4. (3)
    For any $s$ and $\varepsilon$ as in (1), if the group $G(b(s))$ is non-orientable (as a flexible factor), and if the edge group $G([u,b(s)])$ is orientable, then the cardinality of the set of arms $K^{*}_{s^{\prime}}$ with $s^{\prime}\in N_{u}^{T_{*}}$ and with U-type $\varepsilon$ is even, all these arms are ##\##-orientable, and the ##\##-orientations of half of these arms are compatible with one of the orders of the doubleton $\Sigma^{*}_{u}$, while the ##\##-orientations of the other half of these arms are compatible with the opposite order of $\Sigma^{*}_{u}$.
@@
Recall that we denote the underlying tree of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ by $T_{*}$. Let $\Pi^{a}$ be the partition of $T_{*}$ into b-subtrees described as follows. The first kind of b-subtrees in $\Pi^{a}$ are the trivial subtrees $R$ which consist of a single ”core vertex”, i.e. any vertex $r$ with one of the cores (of a V-tree of extended Whitehead graphs) as the constituent space $K_{r}^{*}$. The b-subtrees in $\Pi^{a}$ of the second kind are described as follows. Cut $T_{*}$ along all core vertices (as mentioned above), and along all wedge core vertices (which are white), and consider the b-subtrees $R$ spanned on all black vertices in any of the connected components obtained by cutting $T_{*}$ in this way. Note that for such subtrees $R$ all constituent spaces $K_{r}^{*}$ at the black vertices $r$ in $R$ are arms (in some V-trees of exteneded Whitehead graphs) or wedge arms. For this reason the b-subtrees $R\in\Pi^{a}$ of the latter form will be called arm-subtrees of $T_{*}$. Define the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ as the consolidation of $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ with respect to the partition $\Pi^{a}$, i.e. as $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}:=(\Theta^{V}_{*}[{\mathcal{R}}_{G}])^{\Pi^{a}}$.
@@
We now pass to describing closer those constituent spaces $K_{R}$ of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ which correspond to the arm-subtrees $R$ of $T_{*}$. We will show that, after a suitable interpretation of types, the restriction of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ to any arm-subtree $R$ is isomorphic to a tree system of internally punctured segments of one of the kinds described in Section 12. As a concequence, by referring to Lemma 12.1 or Lemma 12.2, we will identify each of the constituent spaces $K_{R}$ corresponding to arm-subtrees with an appropriate densely decorated punctured circle. More precisely, note first that each arm or wedge arm (viewed as a constituent space in the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$) contains precisely one C-peripheral. Obviously, we will view this arm (or wedge arm) as an internally punctured segment, with its C-peripheral viewed as a special peripheral, and with its E-peripherals viewed as ordinary peripherals. We associate to the special peripheral of any arm (or wedge arm) $A$ the type $\tau$ coinciding with the U-type of $A$ (as described in the previous subsection). If the above type $\tau$ is orientable (as the U-type), we associate to $A$ the orientation which coincides with the ##\##-orientation (as described also in the previous subsection). Observe that, due to Fact 13.1(1), for any arm-subtree $R$, all arms and wedge arms in the restricted tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]|_{R}$ have the same Y-label $y$, which we call therefore the Y-label associated to $R$. It is not hard to observe that the set of U-types of arms and wedge arms appearing in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]|_{R}$ coincides then with the set $U_{y}$ described in the previous subsection. For any such $R$, we take $U_{y}$ as the corresponding set $\mathcal{T}$ of types. It follows then from Fact 13.1(2)(3) that for any arm-subtree $R$ the restricted tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]|_{R}$ is isomorphic, as equipped with the above mentioned types and orientations, to one of the tree systems of internally punctured segments described in Section 12. More precisely, if $y$ is the label associated to $R$, and if the corresponding flexible factor $G(v)$ for any $v\in y$ is orientable, the restricted tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]|_{R}$ with the above described types and orientations is isomorphic to the tree system $\Psi^{o}_{m}(I^{\circ}_{\rm int})$, as described in Subsection 12.1, where $m$ is the cardinality of the subset $U_{y}\subset U$. Similarly, if the corresponding flexible factor $G(v)$ for any $v\in y$ is non-orientable, then the restricted tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]|_{R}$ is isomorphic (via type and orientation respecting isomorphism) with the tree system $\Psi^{n-o}_{m,n}(I^{\circ}_{\rm int})$, as described in Subsection 12.2, where $m$ and $n$ are the cardinalities of the subsets $U_{y,1},U_{y,2}$ consisting of those types from $U_{y}$ which are orientable and nonorientable, respectively. This completes our analysis of the constituent spaces $K_{R}$ corresponding to the arm-subtrees $R$ in the consolidated tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$.
@@
Let ${\mathcal{P}}^{r}_{G}$ be the connecting system described in Example 11.10, and denote the data which form this system as ${\mathcal{P}}^{r}_{G}=(((K_{j},\Omega_{j}),({\mathcal{D}}_{j},d_{j}))_{j\in J},{\mathcal{B}})$. In view of what was said in the previous paragraph, to conclude the proof of Theorem 10.1 we need to show that the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ is modelled on ${\mathcal{P}}^{r}_{G}$. To do this, we will explicitely describe some ${\mathcal{P}}^{r}_{G}$-atlas ${\mathbf{A}}^{a}$ for $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$. Recall first that we have two kinds of constituent spaces in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$. The constituent spaces of the first kind are (the copies of) the cores $C_{z}$ of the V-trees ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{z})$ of extended Whitehead graphs $\widetilde{W}_{z}$ (these graphs are described in Section 10, as part of the structure of ${\mathcal{R}}_{G}$, and in Section 9). By Lemma 9.6(1), each such core $C_{z}$ (and hence each constituent space of $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ which is its copy) is canonically homeomorphic to the space with peripherals $(X_{z},{\mathcal{D}}_{z})$ associated to the corresponding rigid cluster factor $(G(z),{\mathcal{H}}_{z})$ of $G$. The corresponding rigid cluster vertex $z$ of $T^{r}_{G}$ is then in the same $G$-orbit as precisely one of the vertices $v_{j}$ mentioned in Example 11.10, and therefore there is $g\in G$ which maps $z$ to $v_{j}$. This element $g$, in its action on the boundary $\partial G$, maps the space with peripherals $\partial G(z)=X_{z}$ onto the analogous space with peripherals $\partial G(v_{j})=K_{j}$, and we identify $C_{z}$ (via $\partial G(z)$ and $\partial G(v_{j})$) with the corresponding $K_{j}$, using this map (and this identification essentially does not depend on a choice of $g$ as above, since any two such identifications differ by a decoration-preserving homeomorphism). Thus, for each constituent space of the first kind we get in this way its identification with an appropriate space $K_{j}$, and we take it as the corresponding element of the atlas ${\mathbf{A}}^{a}$. The constituent spaces of the second kind in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ are those spaces $K_{R}$ which correspond to the arm-subtrees $R\subset T_{*}$. By the discussion in the previous subsection, each such $K_{R}$ is naturally a densely decorated punctured circle, with types in the set $U_{y}$, where $y\in Y$ is the Y-label associated to $R$. Recalling that $y$ is a $G$-orbit in the set of flexible vertices in $T^{r}_{G}$, we note that the decorated space with peripherals $K_{j}$ with this index $j\in J$ for which $v_{j}\in y$, is also a densely decorated punctured circle. Recalling further that $U_{y}$ is the set of $G$-orbits of the edges of $T^{r}_{G}$ which are adjacent to any fixed vertex $v\in y$, we deduce (by referring to Lemma 11.4(2)(3)) that $K_{R}$ is then isomorphic to $K_{j}$. Moreover, by referring to Lemma 11.4(3), we can (and will) identify $K_{R}$ with the above mentioned $K_{j}$, in an essentially unique way, via a decoration-preserving homeomorphism which sends each peripheral of any type $\varepsilon\in U_{y}$ in $K_{R}$ to a peripheral of type $\tau(\partial G(u^{\prime}))$ in $K_{j}$, where $u^{\prime}$ is any vertex adjacent to $v_{j}$ such that the edge $[u^{\prime},v_{j}]$ belongs to the $G$-orbit $\varepsilon$. Thus, for each constituent space of the second kind in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$, this gives us an identification with an appropriate space $K_{j}$ of ${\mathcal{P}}^{r}_{G}$, and we take this identification as the corresponding element of the atlas ${\mathbf{A}}^{a}$.
@@
It remains to show that the above desribed family ${\mathbf{A}}^{a}$ of identifications satisfies conditions (1)-(4) of Definition 11.8 (as applied to the connecting system ${\mathcal{P}}^{r}_{G}$). Fix the notation $(T^{a},\{K^{a}_{\tilde{t}}\},\{\Sigma^{a}_{\tilde{u}}\},\{\varphi^{a}_{\tilde{e}}=\varphi^{a}_{[\tilde{u},\tilde{t}]}\})$ for the data in $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$. We pass to checking that condition (1) of the Definition 11.8 holds. Let $\Sigma^{a}_{\tilde{u}}$ be some peripheral space in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$. We need to show that the set of types (induced by the above identifications) of the corresponding paripheral subspaces $\varphi^{a}_{[\tilde{u},\tilde{t}]}(\Sigma^{a}_{\tilde{u}})$, where $\tilde{t}$ runs through all vertices adjacent to $\tilde{u}$ in the tree $T^{a}$, coincides with some single block of $\mathcal{B}$. To do this, we consider first the case when $\Sigma^{a}_{\tilde{u}}$ is a core-peripheral (as explained in the previous subsection). Denote by $\tilde{v}$ the (unique) core vertex of $T^{a}$ adjacent to $\tilde{u}$, where by a core vertex we mean here a vertex for which the constituent space $K^{a}_{\tilde{v}}$ is a copy of some core $C_{z}=\partial G(z)$. View then the peripheral $P=\varphi^{a}_{[\tilde{u},\tilde{v}]}(\Sigma^{a}_{\tilde{u}})$ of $K^{a}_{\tilde{v}}$ as a peripheral in the space $\partial G(z)$ (viewed as a space with peripherals associated to the rigid cluster factor $G(z)$), and denote by $u$ the white vertex of $T^{r}_{G}$ for which $P=\partial G(u)$. Denoting by $[u]\in A$ the $G$-orbit of the vertex $u$, we claim that the set of types of the peripherals of form $\varphi^{a}_{[\tilde{t},\tilde{u}]}(\Sigma^{a}_{\tilde{u}})$ coincides with the block $B_{[u]}$ of $\mathcal{B}$. To see this, note that if we interpret the types of the peripherals in ${\mathcal{P}}^{r}_{G}$ as $G$-orbits of edges, as in Observation 11.11, then the above described identifications from the atlas ${\mathbf{A}}^{a}$ simply preserve those types. Moreover, since the peripherals in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ correspond to the C-peripherals of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, we may view the observations from Facts 13.2 and 13.3, after appropriate reinterpretation, as concerning the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$. Our claim follows then fairly directly from Fact 13.2(0). We pass now to considering the remaining case when $\Sigma^{a}_{\tilde{u}}$ is a wedge-peripheral, corresponding to some wedge core. This wedge core is contained in a V-tree of copies of precisely one $\theta$-graph $\Gamma_{z}$ from the connecting system ${\mathcal{R}}_{G}$, and the vertex $z$ is white in $T^{r}_{G}$ and not adjacent to any rigid cluster vertex. We again denote by $[z]\in A$ the $G$-orbit of the vertex $z$, and we observe (in a way similar as in the previous case) that, in view of Fact 13.3(0), the set of types of the peripherals of form $\varphi^{a}_{[\tilde{t},\tilde{u}]}(\Sigma^{a}_{\tilde{u}})$ coincides with the block $B_{[z]}$ of $\mathcal{B}$. This completes the verification of condition (1).

### v2

@@ 13.1. Description of $\Theta^{V}[{\mathcal{R}}_{G}]$ the tree system of V-trees associated to $G$
      moreover, if $A$ is ##\##-non-orientable and $\varepsilon$ is orientable then, after fixing an auxiliary orientation of $A$ and some orientation for $\varepsilon$, the union of all those E-peripherals $P$ of $A$ for which the peripheral $P^{\prime}$ adjacent to $P$ is contained in an arm or wedge arm $A^{\prime}$ of U-type $\varepsilon$, and for which the orientation induced from the fixed orientation of $\varepsilon$ (via the induced orientations of $A^{\prime}$ and $P^{\prime}$) is compatible (respectively, incompatible) with the fixed orientation of $A$, is dense in $A$.
We finish this subsection with another two observations concerning the concepts just introduced, which describe some features of the C-peripherals of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$. These observations follow rather easily from Facts 10.2 and 10.3, respectively, and we omit their proofs.
@@
    Given $s\in N_{u}^{T_{*}}\setminus\{t\}$, consider the subset $[s]\subset N_{u}^{T_{*}}\setminus\{t\}$ corresponding via the bijection $b$ to the $G(u_{\xi})$-orbit $[b(s)]$ of the vertex $b(s)\in N^{T^{r}_{G}}_{u_{\xi}}\setminus\{z\}$. (Equivalently, in accordance with the above part (0), $[s]$ is precisely the set of those $s^{\prime}\in N_{u}^{T_{*}}\setminus\{t\}$ for which the arm $K_{s^{\prime}}^{*}$ has the same U-label as $K^{*}_{s}$.) Suppose also that the group $G(b(s))$ (as well as $G(b(s^{\prime}))$ for any s[$s^{\prime}\in[s]$) is orientable, so that all arms $K^{*}_{s^{\prime}}:s^{\prime}\in[s]$ are ##\##-oriented. Then:
    1. (A)
      if $G(u_{\xi})$ is orientable, then the ##\##-orientations of all arms $K^{*}_{s^{\prime}}:s^{\prime}\in[s]$ are compatible, i.e., they all induce (via the maps $(\varphi_{[s,u]}^{*})^{-1}$) the same order on the doubleton $\Sigma^{*}_{u}$;
    2. (B)
      if $G(u_{\xi})$ is non-orientable, then the cardinality of $[s]$ is even, and the ##\##-orientations of half of the arms $K^{*}_{s^{\prime}}:s^{\prime}\in[s]$ are compatible with one of the orders of the doubleton $\Sigma_{u}^{*}$, while the ##\##-orientations of the other half of these arms are compatible with the opposite order of $\Sigma^{*}_{u}$.
@@
    Given $s\in N_{u}^{T_{*}}\setminus\{t\}$, consider the subset $[s]\subset N_{u}^{T_{*}}$ corresponding via the bijection $b$ to the $G(z)$-orbit $[b(s)]$ of the vertex $b(s)\in N^{T^{r}_{G}}_{z}$. (Equivalently, in accordance with the above part (0), $[s]$ is precisely the set of those $s^{\prime}\in N_{u}^{T_{*}}$ for which the wedge arm $K_{s^{\prime}}^{*}$ has the same U-label as $K^{*}_{s}$.) Suppose also that the group $G(b(s))$ (as well as $G(b(s^{\prime}))$ for any s[$s^{\prime}\in[s]$) is orientable, so that all wedge arms $K^{*}_{s^{\prime}}:s^{\prime}\in[s]$ are ##\##-oriented. Then:
    1. (A)
      if $G(z)$ is orientable, then the ##\##-orientations of all arms $K^{*}_{s^{\prime}}:s^{\prime}\in[s]$ are compatible, i.e., they all induce (via the maps $(\varphi_{[s,u]}^{*})^{-1}$) the same order on the doubleton $\Sigma^{*}_{u}$;
    2. (B)
      if $G(z)$ is non-orientable, then the cardinality of $[s]$ is even, and the ##\##-orientations of half of the arms $K^{*}_{s^{\prime}}:s^{\prime}\in[s]$ are compatible with one of the orders of the doubleton $\Sigma_{u}^{*}$, while the ##\##-orientations of the other half of these arms are compatible with the opposite order of $\Sigma^{*}_{u}$.
@@
Recall that we denote the underlying tree of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ by $T_{*}$. Let $\Pi^{a}$ be the partition of $T_{*}$ into b-subtrees described as follows. The first kind of b-subtrees in $\Pi^{a}$ are the trivial subtrees $R$ which consist of a single “core vertex”, i.e., a vertex $r$ with one of the cores (of a V-tree of extended Whitehead graphs) as the constituent space $K_{r}^{*}$. The b-subtrees in $\Pi^{a}$ of the second kind are described as follows. Cut $T_{*}$ along all core vertices (as mentioned above), and along all wedge core vertices (which are white), and consider the b-subtrees $R$ spanned on all black vertices in any of the connected components obtained by cutting $T_{*}$ in this way. Note that for such subtrees $R$ all constituent spaces $K_{r}^{*}$ at the black vertices $r$ in $R$ are arms (in some V-trees of exteneded Whitehead graphs) or wedge arms. For this reason the b-subtrees $R\in\Pi^{a}$ of the latter form will be called arm-subtrees of $T_{*}$. Define the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ as the consolidation of $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ with respect to the partition $\Pi^{a}$, i.e., as $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}:=(\Theta^{V}_{*}[{\mathcal{R}}_{G}])^{\Pi^{a}}$.
@@
We now pass to describing closer those constituent spaces $K_{R}$ of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ which correspond to the arm-subtrees $R$ of $T_{*}$. We will show that, after a suitable interpretation of types, the restriction of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$ to any arm-subtree $R$ is isomorphic to a tree system of internally punctured segments of one of the kinds described in Section 12. As a concequence, by referring to Lemma 12.1 or Lemma 12.2, we will identify each of the constituent spaces $K_{R}$ corresponding to arm-subtrees with an appropriate densely decorated punctured circle. More precisely, note first that each arm or wedge arm (viewed as a constituent space in the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$) contains precisely one C-peripheral. Obviously, we will view this arm (or wedge arm) as an internally punctured segment, with its C-peripheral viewed as a special peripheral, and with its E-peripherals viewed as ordinary peripherals. We associate to the special peripheral of any arm (or wedge arm) $A$ the type $\tau$ coinciding with the U-type of $A$ (as described in the previous subsection). Moreover, if the Y-label $y$ associated to $A$ (as in the previous subsection) is orientable (i.e. if the flexible factor $G(v)$, for any $v\in y$, is orientable), we associate to $A$ the orientation which coincides with the ##\##-orientation (as described also in the previous subsection). Observe that, due to Fact 13.1(1), for any arm-subtree $R$, all arms and wedge arms in the restricted tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]|_{R}$ have the same Y-label $y$, which we call therefore the Y-label associated to $R$. It is not hard to observe that the set of U-types of arms and wedge arms appearing in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]|_{R}$ coincides then with the set $U_{y}$ described in the previous subsection. For any such $R$, we take $U_{y}$ as the corresponding set $\mathcal{T}$ of types. It follows then from Fact 13.1(2)(3) that for any arm-subtree $R$ the restricted tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]|_{R}$ is isomorphic, as equipped with the above mentioned types and orientations, to one of the tree systems of internally punctured segments described in Section 12. More precisely, if $y$ is the Y-label associated to $R$, and if the corresponding flexible factor $G(v)$ for any $v\in y$ is orientable, the restricted tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]|_{R}$ with the above described types and orientations is isomorphic to the tree system $\Psi^{\text{o}}_{m}(I^{\circ}_{\rm int})$, as described in Section 12.1, where $m$ is the cardinality of the subset $U_{y}\subset U$. Similarly, if the corresponding flexible factor $G(v)$ for any $v\in y$ is non-orientable, then the restricted tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]|_{R}$ is isomorphic (via type and orientation respecting isomorphism) with the tree system $\Psi^{\text{n-o}}_{m,n}(I^{\circ}_{\rm int})$, as described in Section 12.2, where $m$ and $n$ are the cardinalities of the subsets $U_{y,1},U_{y,2}$ consisting of those types from $U_{y}$ which are orientable and nonorientable, respectively. This completes our analysis of the constituent spaces $K_{R}$ corresponding to the arm-subtrees $R$ in the consolidated tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$.
@@
Let ${\mathcal{P}}^{r}_{G}$ be the connecting system described in Example 11.10, and denote the data which form this system as ${\mathcal{P}}^{r}_{G}=(((K_{j},\Omega_{j}),({\mathcal{D}}_{j},d_{j}))_{j\in J},{\mathcal{B}})$. In view of what was said in the previous paragraph, to conclude the proof of Theorem 10.1 we need to show that the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ is modelled on ${\mathcal{P}}^{r}_{G}$. To do this, we will explicitly describe some ${\mathcal{P}}^{r}_{G}$-atlas ${\mathbf{A}}^{a}$ for $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$. Recall first that we have two kinds of constituent spaces in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$. The constituent spaces of the first kind are (the copies of) the cores $C_{z}$ of the V-trees ${\mathcal{X}}_{V}(\widetilde{\mathcal{V}}_{z})$ of extended Whitehead graphs $\widetilde{W}_{z}$ (these graphs are described in Section 10, as part of the structure of ${\mathcal{R}}_{G}$, and in Section 9). By Lemma 9.6(1), each such core $C_{z}$ (and hence each constituent space of $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ which is its copy) is canonically homeomorphic to the space with peripherals $(X_{z},{\mathcal{D}}_{z})$ associated to the corresponding rigid cluster factor $(G(z),{\mathcal{H}}_{z})$ of $G$. The corresponding rigid cluster vertex $z$ of $T^{r}_{G}$ is then in the same $G$-orbit as precisely one of the vertices $v_{j}$ mentioned in Example 11.10, and therefore there is $g\in G$ which maps $z$ to $v_{j}$. This element $g$, in its action on the boundary $\partial G$, maps the space with peripherals $\partial G(z)=X_{z}$ onto the analogous space with peripherals $\partial G(v_{j})=K_{j}$, and we identify $C_{z}$ (via $\partial G(z)$ and $\partial G(v_{j})$) with the corresponding $K_{j}$, using this map (and this identification essentially does not depend on a choice of $g$ as above, since any two such identifications differ by a decoration-preserving homeomorphism). Thus, for each constituent space of the first kind we get in this way its identification with an appropriate space $K_{j}$, and we take it as the corresponding element of the atlas ${\mathbf{A}}^{a}$. The constituent spaces of the second kind in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ are those spaces $K_{R}$ which correspond to the arm-subtrees $R\subset T_{*}$. By the discussion in the previous subsection, each such $K_{R}$ is naturally a densely decorated punctured circle, with types in the set $U_{y}$, where $y\in Y$ is the Y-label associated to $R$. Recalling that $y$ is a $G$-orbit in the set of flexible vertices in $T^{r}_{G}$, we note that the decorated space with peripherals $K_{j}$ with this index $j\in J$ for which $v_{j}\in y$, is also a densely decorated punctured circle. Recalling further that $U_{y}$ is the set of $G$-orbits of the edges of $T^{r}_{G}$ which are adjacent to any fixed vertex $v\in y$, we deduce (by referring to Lemma 11.4(2)(3)) that $K_{R}$ is then isomorphic to $K_{j}$. Moreover, by referring to Lemma 11.4(2)(3), we can (and will) identify $K_{R}$ with the above mentioned $K_{j}$, in an essentially unique way, via a decoration-preserving homeomorphism which sends each peripheral of any type $\varepsilon\in U_{y}$ in $K_{R}$ to a peripheral of type $\tau(\partial G(u^{\prime}))$ in $K_{j}$, where $u^{\prime}$ is any vertex adjacent to $v_{j}$ such that the edge $[u^{\prime},v_{j}]$ belongs to the $G$-orbit $\varepsilon$. Thus, for each constituent space of the second kind in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$, this gives us an identification with an appropriate space $K_{j}$ of ${\mathcal{P}}^{r}_{G}$, and we take this identification as the corresponding element of the atlas ${\mathbf{A}}^{a}$.
@@
It remains to show that the above described family ${\mathbf{A}}^{a}$ of identifications satisfies conditions (1)-(4) of Definition 11.8 (as applied to the connecting system ${\mathcal{P}}^{r}_{G}$). Fix the notation $(T^{a},\{K^{a}_{\tilde{t}}\},\{\Sigma^{a}_{\tilde{u}}\},\{\varphi^{a}_{\tilde{e}}=\varphi^{a}_{[\tilde{u},\tilde{t}]}\})$ for the data in $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$. We pass to checking that condition (1) of the Definition 11.8 holds. Let $\Sigma^{a}_{\tilde{u}}$ be some peripheral space in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$. We need to show that the set of types (induced by the above identifications) of the corresponding peripheral subspaces $\varphi^{a}_{[\tilde{u},\tilde{t}]}(\Sigma^{a}_{\tilde{u}})$, where $\tilde{t}$ runs through all vertices adjacent to $\tilde{u}$ in the tree $T^{a}$, coincides with some single block of $\mathcal{B}$. To do this, we consider first the case when $\Sigma^{a}_{\tilde{u}}$ is a core-peripheral (as explained in the previous subsection). Denote by $\tilde{v}$ the (unique) core vertex of $T^{a}$ adjacent to $\tilde{u}$, where by a core vertex we mean here a vertex for which the constituent space $K^{a}_{\tilde{v}}$ is a copy of some core $C_{z}=\partial G(z)$. View then the peripheral $P=\varphi^{a}_{[\tilde{u},\tilde{v}]}(\Sigma^{a}_{\tilde{u}})$ of $K^{a}_{\tilde{v}}$ as a peripheral in the space $\partial G(z)$ (viewed as a space with peripherals associated to the rigid cluster factor $G(z)$), and denote by $u$ the white vertex of $T^{r}_{G}$ for which $P=\partial G(u)$. Denoting by $[u]\in A$ the $G$-orbit of the vertex $u$ (where $A$ is as in Example 11.10), we claim that the set of types of the peripherals of the form $\varphi^{a}_{[\tilde{t},\tilde{u}]}(\Sigma^{a}_{\tilde{u}})$ coincides with the block $B_{[u]}$ of $\mathcal{B}$. To see this, note that if we interpret the types of the peripherals in ${\mathcal{P}}^{r}_{G}$ as $G$-orbits of edges, as in Observation 11.11, then the above described identifications from the atlas ${\mathbf{A}}^{a}$ simply preserve those types. Moreover, since the peripherals in the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$ correspond to the C-peripherals of the system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]$, we may view the observations from Facts 13.2 and 13.3, after appropriate reinterpretation, as concerning the tree system $\Theta^{V}_{*}[{\mathcal{R}}_{G}]^{a}$. Our claim follows then fairly directly from Fact 13.2(0). We pass now to considering the remaining case when $\Sigma^{a}_{\tilde{u}}$ is a wedge-peripheral, corresponding to some wedge core. This wedge core is contained in a V-tree of copies of precisely one $\theta$-graph $\Gamma_{z}$ from the connecting system ${\mathcal{R}}_{G}$, and the vertex $z$ is white in $T^{r}_{G}$ and not adjacent to any rigid cluster vertex. We denote by $[z]\in A$ the $G$-orbit of the vertex $z$, and we observe (in a way similar as in the previous case) that, in view of Fact 13.3(0), the set of types of the peripherals of the form $\varphi^{a}_{[\tilde{t},\tilde{u}]}(\Sigma^{a}_{\tilde{u}})$ coincides with the block $B_{[z]}$ of $\mathcal{B}$. This completes the verification of condition (1).